# World Model on Million-Length Video And Language With Blockwise RingAttention

### v2

@@ 1 Introduction
Current approaches on modeling the world are mostly restricted to short sequences of language sequences or short sequences of images and clips  (${}^{+}$, 20; ${}^{+}$, 23; ${}^{+}$, 23; Ope, 23; ${}^{+}$, 23). resulting in models which lack understanding about parts of the world that are hard to represent in texts or short clips, and are unable to process complex long-form language and visual tasks. Temporal structure in video sequences provides helpful information that is missing from language or far less obvious in static images and short clips. Long language sequences encode information that short sequences cannot, crucial for various applications such as long document retrieval or coding. Long videos provide a rich context that short clips cannot grasp, showing how scenes connect, the development of events, and the cause and effect of actions within the temporal dimension of the video. This exposure to diverse long language and video scenarios also broadens the AI systems to generalize across various real-world situations. By jointly modeling both long videos and books, the model can develop an understanding of both the multimodal world and long sequences of texts and videos, leading to more advanced AI systems with a multimodal understanding, capable of assisting humans in a broader range of tasks.
@@
To learn from video and language sequences, we need to train a model that is capable of processing more than millions of tokens per sequence and train it on a very large dataset. However, modeling millions of tokens is extremely difficult due to high memory cost, computational complexity, and lack of suitable datasets. Luckily, we have Blockwise RingAttention (LZA, 24; LA, 23), a technique for scaling up context size arbitrarily without approximations or overheads, allowing for scalable training on long sequences. We curated a large dataset of videos and languages from public book and video datasets, consisting of videos of diverse activities and long-form books. Considering the high compute cost, we grow context size from a smaller 4K context to a larger 1M context size gradually to reduce this cost, and this approach performs well in extending context effectively. Furthermore, we identify challenges associated with training on video and language: we discovered that training on a mixture of video, image, and text is crucial for optimal performance, due to images represent higher visual quality, videos offer sequential information, and text retains language understanding. To achieve this, we implemented an efficient masked sequence packing to effectively train with different sequence lengths, rather than standard sequence packing mechanism. Moreover, determining the right balance between image, video, and text training is crucial for cross modality understanding, and we suggest a ratio that proved effective. Furthermore, to address the lack of long-form chat datasets, we developed a model-generated question-answering (QA) approach by using a short-context model to generate a QA dataset from books. We found this to be crucial for chat abilities over long sequences.
@@ 2 Overview
We train a large autoregressive transformer model with a very large context window of up to one million tokens, building upon Llama2 7B (${}^{+}$, 23). To achieve this goal, we leverage several strategies: extending the context to 1M using books (Section 3), followed by joint training on long multimodal sequences, including text-image, text-video data, and books (Section 4).
@@ 3.1 Extending Context
Scalable Training on Long Documents. Training on long documents becomes prohibitively expensive due to memory constraints imposed by the quadratic complexity of computing the attention weights. In order to address these computational constraints, we use the Blockwise RingAttention (LZA, 24; LA, 23) implementation that leverages block-wise transformer with sequence parallelism to theoretically extend to an infinite context, bounded only by the number of devices available. We further fuse Blockwise RingAttention with FlashAttention (${}^{+}$, 22; RS, 21) using Pallas (${}^{+}$, 18) to optimize performance compared with using XLA compiler. In general, given a large enough tokens per device, the communication cost during Blockwise Transformer and RingAttention fully overlap with computation, and does not add any extra overhead.
@@
Therefore, we adopt a training approach inspired from ${}^{+}$ (23), where our model is trained on progressively longer sequence lengths, starting from 32K tokens and ending at 1M tokens in increasing powers of two. Intuitively, this allows the model to save compute by first learning shorter-range dependencies before moving onto longer sequences. By doing this, we are able to train on orders of magnitude more tokens compared to directly training on the maximum target sequence length. The progressive training of growing context sizes is shown in Figure 3.
Positional Extrapolation for Long Contexts. For extending positional embeddings on longer contexts, we adopt a simple, scaled up version of the approach explored in ${}^{+}$ (23), where the $\theta$ for RoPE (${}^{+}$, 24) is scaled up with context length. We generally found this approach to be a stable method to extend positional embeddings with context lengths due to its relatively simple nature of only needing to tune a single hyperparameter. We scale up the $\theta$ for RoPE along with context window sizes – the values are shown in Table 1.
@@ 3.2 Training Steps
We initialize from LLaMA-2 7B ${}^{+}$ (23) and progressively increase the effective context length of the model in 5 stages - 32K, 128K, 256K, 512K, and 1M. For each stage, we train on different filtered versions of the Books3 dataset from The Pile ${}^{+}$ (20). Table 1 details each information about each training stage, such as number of tokens, total time, and the Books3 dataset filtering constraints. Each successive run is initialized from the run of the prior sequence length.
@@ 3.3 Chat Fine-tuning for Long-Context Learning
Training Details.  For chat fine-tuning, we train each model on a mix of UltraChat ${}^{+}$ (23) and our custom QA dataset, with roughly a 7:3 ratio. We found it crucial to pre-pack the UltraChat data to the training sequence length, and keep them separate from examples with our QA data, as UltraChat data generally has a much higher proportion of loss tokens (densely packed, short chat sequences), whereas our QA data has a much lower percentage of loss tokens per sequence (< 1%) since there is no loss on the long documents that are in the given context. Table 2 shows further training details for each run. Note that progressive training is not performed very each of the chat models, and instead are initialized from their respective pretrained models at the same context length.
@@ 3.4.1 Single Needle Retrieval
We evaluate on the popular Needle In A Haystack task (gka, 23) - more specifically an easier to evaluate version (AI, 23) that finds and retrieves random numbers assigned to randomized cities from the context. Figure 5 shows nearly perfect retrieval accuracy over the entire context of our 1M context model. In addition, Figure 2 shows that we can scale to far larger contexts compared to the current best available LLMs. Appendix A shows more single needle retrieval results for our other shorter context length models.
@@ 3.4.4 Chat Evaluation
We additionally evaluate the our model on MT-Bench (${}^{+}$23a, ) to test its conversation ability. Table 6 shows the MT-Bench scores of for each of our models. Table 6 illustrates the relationship between the mix of chat and fact retrieval tasks and the performance on MT-Bench score and Needle Retrieval accuracy. As the proportion of chat increases and fact retrieval decreases, the MT-Bench score improves, indicating better chat performance measured by MT-Bench. Conversely, Needle Retrieval accuracy decreases, suggesting a trade-off where increasing chat interaction capabilities may reduce the system’s precision in retrieving specific information or ’needles’ from input context. Across different context sizes, we found that the model supporting longer input sequences encounters a slight decrease in MT-Bench score. We hypothesize that this is because we chose to train with fewer examples on longer sequence training and can be improved by simply training on more data. In addition, this trade-off may be resolved by acquiring higher quality long-context chat data that is closer to the chat distribution of the UltraChat dataset.
@@ 4.1 Architectural Modifications For Vision
The model is illustrated in Figure 4. We use the pretrained VQGAN ERO (21) from aMUSEd PBRvP (24) that tokenizes $256\times 256$ input images to $16\times 16$ discrete tokens. Videos are tokenized by applying the VQGAN per-frame, and concatenating the codes together. In order to distinguish between modalities when generating, as well as knowing when to switch, we introduce mechanims to mark the end of text generation / beginning of vision generation, and vice-versa. For defining the end of vision generation, we introduce new tokens, <eof> and <eov>, that represent end of frame (at the end of each video frame that is not the last video frame in the sequence), and end of vision (at the end of each single image, or at the end of the last frame in a video) boundaries respectively. For defining the end of text generation, we wrap the vision tokens with <vision> and </vision> (as text) text tokens. The model is trained with interleaved concatenations of vision and text tokens, and predicted autoregressively.
@@ 4.2 Training Steps
  • LWM-1K: We train on large set of text-image dataset comprising of a mix of LAION-2B-en ${}^{+}$ (22) and COYO-700M ${}^{+}$ (22). The datasets were filtered to only include images with at least 256 resolution – in total roughly 1B text-image pairs. During training, we concatenate the text-image pairs and randomly swap the order of the modalities to model both text-image generation, unconditional image generation, and image captioning. We pack text-image pairs to sequences of 1K tokens.
  • LWM-8K: We train on a text-video dataset mix of WebVid10M BNVZ (21) and 3M InternVid10M ${}^{+}$ (23) examples. Similar to prior works ${}^{+}$ (22); ${}^{+}$ (22); ${}^{+}$ (22), we jointly train on both images and video with a 50-50 ratio of each modality. We pack images to sequences of 8K tokens, and 30 frame videos at 4FPS. Similar to image training, we randomly swap the order of modalities for each text-video pair.
  • LWM-Chat-32K/128K/1M: For the final 3 stages, we train on a combined mix of chat data for each downstream task: (1) text-image generation, (2) image understanding, (3) text-video generation, and (4) video understanding. We construct a simple version of text-image and text-video chat data by sampling random subsets of the pretraining data augmented with chat format. For image understanding, we using the image chat instruct data from ShareGPT4V ${}^{+}$ (23). Lastly, for the video understanding chat data, we use a combined mix of Valley-Instruct-73K ${}^{+}$23b and Video-ChatGPT-100K instruct data MRKK (23). For all short context data (image generation, image understanding, video generation), we pack sequences to the training context length. During packing, we found it crucial to mask out the attention so that each text-vision pair only attends to itself, as well as re-weighting losses to make computation identical to training in a non-packed + padding training regime. For video understanding data, we uniformly sample a max number of frames to fit the training context length of the model if the video is too long. During training, We allocate 25% of each batch to each of the 4 downstream tasks.
For the first two stages of training (LWM-1K and LWM-8K), we additionally mix 16% of the batch to be pure text data from OpenLLaMA GL (23), as we found it beneficial to preserve language capabilities while training on vision data.
@@ 4.3.1 Long Video Understanding
Although vision-language model ${}^{+}$23a ; Ope (23); ${}^{+}$ (23) can ingest long videos, this is commonly done by performing large temporal subsampling of video frames due to limited context length. For example, Video-LLaVA ${}^{+}$23a is restricted to uniformly sampling 8 frames from a video, no matter how long the original video may be. As such, models may lose more fine-grained temporal information that is important for accurately answering any questions about the video. In contrast, our model is trained on long sequences of 1M tokens, and as a result, can simultaneously attend thousands of frames of videos to retrieve fine-grained information over short time intervals. Figure 1 shows an example of our model correctly answering questions about a long, 1-hour YouTube compilation consisting of more than 500 individual clips. Our baseline methods, on the other hand, generally have difficulty answering the questions due to a limited number of frames. More results are shown in Figure 7 and Appendix B.
@@ 4.3.2 Image Understanding and Short Video Understanding
Method Visual Token VQAv2 GQA VisWiz SQA TextVQA POPE MM-Vet
MiniGPT-4 ${}^{+}$23b CLIP - 30.8 47.5 25.4 19.4 - 22.1
Otter ${}^{+}$ (23) CLIP - 38.1 50 27.2 21.2 - 24.6
InstructBLIP ${}^{+}$ (23) CLIP - 49.2 34.5 60.5 50.1 - 26.2
LLaVA-1.5 LLLL (23) CLIP 78.5 62 38.9 66.8 58.2 85.9 30.5
LWM (ours) VQGAN 55.8 44.8 11.6 47.7 18.8 75.2 9.6
@@
MSVD-QA MSRVTT-QA TGIF-QA
Method Visual Token Accuracy Score Accuracy Score Accuracy Score
VideoChat ${}^{+}$ (23) CLIP 56.3 2.8 45 2.5 34.4 2.3
LLaMA-Adapter ${}^{+}$ (23) CLIP 54.9 3.1 43.8 2.5 - -
Video-LLaMA ZLB (23) CLIP 51.6 2.5 29.6 1.8 - -
Video-ChatGPT MRKK (23) CLIP 64.9 3.3 49.3 2.8 51.4 3
Video-LLaVA ${}^{+}$23a CLIP 70.7 3.9 59.2 3.5 70 4
LWM (ours) VQGAN 55.9 3.5 44.1 3.1 40.9 3.1
@@ 4.3.3 Image and Video Generation
In addition to being able to perform image / video captions, as well as QA, our model can also generate images and video from text. Figure 8 shows some examples of such capabilities. We use classifier-free guidance HS (22) on the logits for autoregressive sampling in a manner similar to prior work ${}^{+}$ (22); ${}^{+}$ (22). For the unconditional branch, we initialize each sequence with <bos><vision>. Appendix E and D show more image and video generation examples.
@@ 5 Further Details
MFU.  We trained our models using TPUv4-1024, which is approximately equivalent to 450 A100s, with a batch size of 8M using FSDP Fac (23), Blockwise Transformer, and RingAttention for large contexts. Figure 9 shows the model FLOPS utilization (MFU) for each training stage. Blue color bars show language training and orange color bars show vision-language training. Our training achieves good MFUs even for very large context sizes.
@@ 6 Related Works
Our work is related to efforts that extend language models’ context windows to allow more tokens (CWCT, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia), often by using novel extrapolation methods to expand pretrained positional encodings followed by finetuning the model on longer context data. Our model adopts a simple approach by gradually increasing $\theta$ in RoPE positional encodings along with the training context window sizes, and we found this method to be effective. There has been research into architectures that do not model pairwise interaction, such as sparse attention and sliding window (CGRS, 19; BPC, 20). Parallelization along the sequence dimension is studied in (LXLY, 21), but is not optimized for blockwise transformers and is incompatible with memory-efficient attention, which are crucial to large context training. (${}^{+}$, 22) propose hybrid usage of tensor parallelism and sequence parallelism to better reduce memory cost on a single device. Our work utilizes RingAttention and Blockwise Transformer (LZA, 24; LA, 23) to model exact pairwise interactions in very long sequences for optimal performance. Further training performance improvement is also possible with the load balancing of skipping causal masked computation (${}^{+}$, 23; ${}^{+}$, 23).
Our work is also related to works on instruction tuning (${}^{+}$, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia). These studies focus on finetuning models using conversational data to enhance their capability across language tasks. Our approach seeks to advance models’ understanding of complex, long sequences of videos and languages. To achieve this, we extend the models’ context size by training on books and long videos, and finetuning on model-generated QA data to learn chat ability over long sequences.
Our work is also related to efforts combining vision and language (LLWL, 23; ${}^{+}$23a, ; ${}^{+}$, 23; ${}^{+}$, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia). These efforts often use CLIP (${}^{+}$, 21) or BLIP (LLXH, 22) to encode visual information into embeddings for inputting into language models. They have the potential advantages of leveraging CLIP’s cross-modal understanding for encoding textual information in images. However, they can only be trained to predict text given visual input but not vice versa, limiting their ability to learn from diverse formats of visual and language information. In cases where they do predict visual tokens ${}^{+}$ (23), a stronger diffusion decoder is generally required due to the relatively lossy nature of CLIP embeddings. Our work, on the other hand, is autoregressive "tokens in, tokens out", allowing us to flexibly model diverse forms of image-text, text-image, text-video, video-text, and purely video, image, or text formats.

### v3

@@ 1 Introduction
Current approaches on modeling the world are mostly restricted to short sequences of language sequences or short sequences of images and clips  [5, 54, 55, 44, 52]. resulting in models which lack understanding about parts of the world that are hard to represent in texts or short clips, and are unable to process complex long-form language and visual tasks. Temporal structure in video sequences provides helpful information that is missing from language or far less obvious in static images and short clips. Long language sequences encode information that short sequences cannot, crucial for various applications such as long document retrieval or coding. Long videos provide a rich context that short clips cannot grasp, showing how scenes connect, the development of events, and the cause and effect of actions within the temporal dimension of the video. This exposure to diverse long language and video scenarios also broadens the AI systems to generalize across various real-world situations. By jointly modeling both long videos and books, the model can develop an understanding of both the multimodal world and long sequences of texts and videos, leading to more advanced AI systems with a multimodal understanding, capable of assisting humans in a broader range of tasks.
@@
To learn from video and language sequences, we need to train a model that is capable of processing more than millions of tokens per sequence and train it on a very large dataset. However, modeling millions of tokens is extremely difficult due to high memory cost, computational complexity, and lack of suitable datasets. Luckily, we have Blockwise RingAttention [39, 31], a technique for scaling up context size arbitrarily without approximations or overheads, allowing for scalable training on long sequences. We curated a large dataset of videos and languages from public book and video datasets, consisting of videos of diverse activities and long-form books. Considering the high compute cost, we grow context size from a smaller 4K context to a larger 1M context size gradually to reduce this cost, and this approach performs well in extending context effectively. Furthermore, we identify challenges associated with training on video and language: we discovered that training on a mixture of video, image, and text is crucial for optimal performance, due to images represent higher visual quality, videos offer sequential information, and text retains language understanding. To achieve this, we implemented an efficient masked sequence packing to effectively train with different sequence lengths, rather than standard sequence packing mechanism. Moreover, determining the right balance between image, video, and text training is crucial for cross modality understanding, and we suggest a ratio that proved effective. Furthermore, to address the lack of long-form chat datasets, we developed a model-generated question-answering (QA) approach by using a short-context model to generate a QA dataset from books. We found this to be crucial for chat abilities over long sequences.
@@ 2 Overview
We train a large autoregressive transformer model with a very large context window of up to one million tokens, building upon Llama2 7B [55]. To achieve this goal, we leverage several strategies: extending the context to 1M using books (Section 3), followed by joint training on long multimodal sequences, including text-image, text-video data, and books (Section 4).
@@ 3.1 Extending Context
Scalable Training on Long Documents. Training on long documents becomes prohibitively expensive due to memory constraints imposed by the quadratic complexity of computing the attention weights. In order to address these computational constraints, we use the Blockwise RingAttention [39, 31] implementation that leverages block-wise transformer with sequence parallelism to theoretically extend to an infinite context, bounded only by the number of devices available. We further fuse Blockwise RingAttention with FlashAttention [15, 48] using Pallas [4] to optimize performance compared with using XLA compiler. In general, given a large enough tokens per device, the communication cost during Blockwise Transformer and RingAttention fully overlap with computation, and does not add any extra overhead.
@@
Therefore, we adopt a training approach inspired from [28], where our model is trained on progressively longer sequence lengths, starting from 32K tokens and ending at 1M tokens in increasing powers of two. Intuitively, this allows the model to save compute by first learning shorter-range dependencies before moving onto longer sequences. By doing this, we are able to train on orders of magnitude more tokens compared to directly training on the maximum target sequence length. The progressive training of growing context sizes is shown in Figure 3.
Positional Extrapolation for Long Contexts. For extending positional embeddings on longer contexts, we adopt a simple, scaled up version of the approach explored in [46], where the $\theta$ for RoPE [50] is scaled up with context length. We generally found this approach to be a stable method to extend positional embeddings with context lengths due to its relatively simple nature of only needing to tune a single hyperparameter. We scale up the $\theta$ for RoPE along with context window sizes – the values are shown in Table 1.
@@ 3.2 Training Steps
We initialize from LLaMA-2 7B [55] and progressively increase the effective context length of the model in 5 stages - 32K, 128K, 256K, 512K, and 1M. For each stage, we train on different filtered versions of the Books3 dataset from The Pile [19]. Table 1 details each information about each training stage, such as number of tokens, total time, and the Books3 dataset filtering constraints. Each successive run is initialized from the run of the prior sequence length.
@@ 3.3 Chat Fine-tuning for Long-Context Learning
Training Details.  For chat fine-tuning, we train each model on a mix of UltraChat [14] and our custom QA dataset, with roughly a 7:3 ratio. We found it crucial to pre-pack the UltraChat data to the training sequence length, and keep them separate from examples with our QA data, as UltraChat data generally has a much higher proportion of loss tokens (densely packed, short chat sequences), whereas our QA data has a much lower percentage of loss tokens per sequence (< 1%) since there is no loss on the long documents that are in the given context. Table 2 shows further training details for each run. Note that progressive training is not performed very each of the chat models, and instead are initialized from their respective pretrained models at the same context length.
@@ 3.4.1 Single Needle Retrieval
We evaluate on the popular Needle In A Haystack task [22] - more specifically an easier to evaluate version [2] that finds and retrieves random numbers assigned to randomized cities from the context. Figure 5 shows nearly perfect retrieval accuracy over the entire context of our 1M context model. In addition, Figure 2 shows that we can scale to far larger contexts compared to the current best available LLMs. Appendix A shows more single needle retrieval results for our other shorter context length models.
@@ 3.4.4 Chat Evaluation
We additionally evaluate the our model on MT-Bench [60] to test its conversation ability. Table 6 shows the MT-Bench scores of for each of our models. Table 6 illustrates the relationship between the mix of chat and fact retrieval tasks and the performance on MT-Bench score and Needle Retrieval accuracy. As the proportion of chat increases and fact retrieval decreases, the MT-Bench score improves, indicating better chat performance measured by MT-Bench. Conversely, Needle Retrieval accuracy decreases, suggesting a trade-off where increasing chat interaction capabilities may reduce the system’s precision in retrieving specific information or ’needles’ from input context. Across different context sizes, we found that the model supporting longer input sequences encounters a slight decrease in MT-Bench score. We hypothesize that this is because we chose to train with fewer examples on longer sequence training and can be improved by simply training on more data. In addition, this trade-off may be resolved by acquiring higher quality long-context chat data that is closer to the chat distribution of the UltraChat dataset.
@@ 4.1 Architectural Modifications For Vision
The model is illustrated in Figure 4. We use the pretrained VQGAN [17] from aMUSEd [45] that tokenizes $256\times 256$ input images to $16\times 16$ discrete tokens. Videos are tokenized by applying the VQGAN per-frame, and concatenating the codes together. In order to distinguish between modalities when generating, as well as knowing when to switch, we introduce mechanims to mark the end of text generation / beginning of vision generation, and vice-versa. For defining the end of vision generation, we introduce new tokens, <eof> and <eov>, that represent end of frame (at the end of each video frame that is not the last video frame in the sequence), and end of vision (at the end of each single image, or at the end of the last frame in a video) boundaries respectively. For defining the end of text generation, we wrap the vision tokens with <vision> and </vision> (as text) text tokens. The model is trained with interleaved concatenations of vision and text tokens, and predicted autoregressively.
@@ 4.2 Training Steps
  • LWM-1K: We train on large set of text-image dataset comprising of a mix of LAION-2B-en [51] and COYO-700M [9]. The datasets were filtered to only include images with at least 256 resolution – in total roughly 1B text-image pairs. During training, we concatenate the text-image pairs and randomly swap the order of the modalities to model both text-image generation, unconditional image generation, and image captioning. We pack text-image pairs to sequences of 1K tokens.
  • LWM-8K: We train on a text-video dataset mix of WebVid10M [7] and 3M InternVid10M [58] examples. Similar to prior works [25, 27, 57], we jointly train on both images and video with a 50-50 ratio of each modality. We pack images to sequences of 8K tokens, and 30 frame videos at 4FPS. Similar to image training, we randomly swap the order of modalities for each text-video pair.
  • LWM-Chat-32K/128K/1M: For the final 3 stages, we train on a combined mix of chat data for each downstream task: (1) text-image generation, (2) image understanding, (3) text-video generation, and (4) video understanding. We construct a simple version of text-image and text-video chat data by sampling random subsets of the pretraining data augmented with chat format. For image understanding, we using the image chat instruct data from ShareGPT4V [11]. Lastly, for the video understanding chat data, we use a combined mix of Valley-Instruct-73K [42] and Video-ChatGPT-100K instruct data [43]. For all short context data (image generation, image understanding, video generation), we pack sequences to the training context length. During packing, we found it crucial to mask out the attention so that each text-vision pair only attends to itself, as well as re-weighting losses to make computation identical to training in a non-packed + padding training regime. For video understanding data, we uniformly sample a max number of frames to fit the training context length of the model if the video is too long. During training, We allocate 25% of each batch to each of the 4 downstream tasks.
For the first two stages of training (LWM-1K and LWM-8K), we additionally mix 16% of the batch to be pure text data from OpenLLaMA [23], as we found it beneficial to preserve language capabilities while training on vision data.
@@ 4.3.1 Long Video Understanding
Although vision-language model [41, 44, 52] can ingest long videos, this is commonly done by performing large temporal subsampling of video frames due to limited context length. For example, Video-LLaVA [41] is restricted to uniformly sampling 8 frames from a video, no matter how long the original video may be. As such, models may lose more fine-grained temporal information that is important for accurately answering any questions about the video. In contrast, our model is trained on long sequences of 1M tokens, and as a result, can simultaneously attend thousands of frames of videos to retrieve fine-grained information over short time intervals. Figure 1 shows an example of our model correctly answering questions about a long, 1-hour YouTube compilation consisting of more than 500 individual clips. Our baseline methods, on the other hand, generally have difficulty answering the questions due to a limited number of frames. More results are shown in Figure 7 and Appendix B.
@@ 4.3.2 Image Understanding and Short Video Understanding
{adjustwidth}
-0.2cm Method Visual Token VQAv2 GQA VisWiz SQA TextVQA POPE MM-Vet MiniGPT-4 [61] CLIP - 30.8 47.5 25.4 19.4 - 22.1 Otter [40] CLIP - 38.1 50 27.2 21.2 - 24.6 InstructBLIP [16] CLIP - 49.2 34.5 60.5 50.1 - 26.2 LLaVA-1.5 [33] CLIP 78.5 62 38.9 66.8 58.2 85.9 30.5 LWM (ours) VQGAN 55.8 44.8 11.6 47.7 18.8 75.2 9.6
@@
{adjustwidth}
-0.2cm MSVD-QA MSRVTT-QA TGIF-QA Method Visual Token Accuracy Score Accuracy Score Accuracy Score VideoChat [32] CLIP 56.3 2.8 45 2.5 34.4 2.3 LLaMA-Adapter [21] CLIP 54.9 3.1 43.8 2.5 - - Video-LLaMA [63] CLIP 51.6 2.5 29.6 1.8 - - Video-ChatGPT [43] CLIP 64.9 3.3 49.3 2.8 51.4 3 Video-LLaVA [41] CLIP 70.7 3.9 59.2 3.5 70 4 LWM (ours) VQGAN 55.9 3.5 44.1 3.1 40.9 3.1
@@ 4.3.3 Image and Video Generation
In addition to being able to perform image / video captions, as well as QA, our model can also generate images and video from text. Figure 8 shows some examples of such capabilities. We use classifier-free guidance [26] on the logits for autoregressive sampling in a manner similar to prior work [59, 24]. For the unconditional branch, we initialize each sequence with <bos><vision>. Appendix E and D show more image and video generation examples.
@@ 5 Further Details
MFU.  We trained our models using TPUv4-1024, which is approximately equivalent to 450 A100s, with a batch size of 8M using FSDP [18], Blockwise Transformer, and RingAttention for large contexts. Figure 9 shows the model FLOPS utilization (MFU) for each training stage. Blue color bars show language training and orange color bars show vision-language training. Our training achieves good MFUs even for very large context sizes.
@@ 6 Related Works
Our work is related to efforts that extend language models’ context windows to allow more tokens [13, 56, 38, inter alia], often by using novel extrapolation methods to expand pretrained positional encodings followed by finetuning the model on longer context data. Our model adopts a simple approach by gradually increasing $\theta$ in RoPE positional encodings along with the training context window sizes, and we found this method to be effective. There has been research into architectures that do not model pairwise interaction, such as sparse attention and sliding window [10, 8]. Parallelization along the sequence dimension is studied in [37], but is not optimized for blockwise transformers and is incompatible with memory-efficient attention, which are crucial to large context training. [30] propose hybrid usage of tensor parallelism and sequence parallelism to better reduce memory cost on a single device. Our work utilizes RingAttention and Blockwise Transformer [39, 31] to model exact pairwise interactions in very long sequences for optimal performance. Concurrent to our work, Gemini 1.5 [49] achieves millions of tokens context size. Further training performance improvement is also possible with the load balancing of skipping causal masked computation [6, 36].
Our work is also related to works on instruction tuning [53, 12, 20, inter alia]. These studies focus on finetuning models using conversational data to enhance their capability across language tasks. Our approach seeks to advance models’ understanding of complex, long sequences of videos and languages. To achieve this, we extend the models’ context size by training on books and long videos, and finetuning on model-generated QA data to learn chat ability over long sequences.
Our work is also related to efforts combining vision and language [34, 41, 1, 62, 29, 3, inter alia]. These efforts often use CLIP [47] or BLIP [35] to encode visual information into embeddings for inputting into language models. They have the potential advantages of leveraging CLIP’s cross-modal understanding for encoding textual information in images. However, they can only be trained to predict text given visual input but not vice versa, limiting their ability to learn from diverse formats of visual and language information. In cases where they do predict visual tokens [29], a stronger diffusion decoder is generally required due to the relatively lossy nature of CLIP embeddings. Our work, on the other hand, is autoregressive "tokens in, tokens out", allowing us to flexibly model diverse forms of image-text, text-image, text-video, video-text, and purely video, image, or text formats.